On March 31, 2021, management of Quality Appliances committed to a plan to sell equipment. The equipment was available for immediate sale, and an active plan to locate a buyer was initiated. The equipment had been purchased on January 1, 2019, for $260,000. The equipment had an estimated six-year service life and residual value of $20,000. The equipment was being depreciated using the straight-line method. Quality’s fiscal year ends on December 31. Required:

Calculate the equipment’s book value as of March 31, 2021 (Hint: Depreciation for 2021 would include up to March 31).

By December 31, 2021, the equipment has not been sold, but management expects that it will be sold in 2022 for $150,000. For what amount is the equipment reported in the December 31, 2021, balance sheet?

Step by step
Answer #1

Cost of equipment = $260,000

Useful life = 6 years

Residual value = $20,000

Annual depreciation = ( Cost of equipment-Residual value)/Useful life

= (260,000-20,000)/6

= 240,000/6

= $40,000

Depreciation expense for year 2019 = $40,000

Depreciation expense for year 2020 = $40,000

Depreciation expense for year 2021 for 3 months (January 1 to March 31)= 40,000 x 3/12

= $10,000

Accumulated depreciation up to March 31, 2021 = Depreciation expense for year 2019 + Depreciation expense for year 2020+ Depreciation expense for year 2021

= 40,000+40,000+10,000

= $90,000

The equipment’s book value as of March 31, 2021 = Cost of equipment - Accumulated depreciation up to March 31, 2021

= 260,000-90,000

= $170,000

Accumulated depreciation on December 31, 2021 = Annual depreciation expense x 3

= 40,000 x 3

= $120,000

Book value of equipment on December 31, 2021 = Cost of equipment - Accumulated depreciation on December 31, 2021

= 260,000-120,000

= $140,000

Amount is the equipment reported in the December 31, 2021, balance sheet $140,000.

Please give a positive rating if you are satisfied with this solution and if you have any query kindly ask.Thanks!!!